OpenStudy (anonymous):

solve each equation. state the number and type of roots. x^4-81=0

OpenStudy (cwrw238):

adding 81 to each side x^4 = 81 now you need to find x where x * x * x * x = 81

OpenStudy (cwrw238):

hint: 9 * 9 = 81

OpenStudy (anonymous):

i figured that x^4= 81 adnd then im lost

OpenStudy (cwrw238):

and x * x = 9

OpenStudy (cwrw238):

so whats x equal to ?

OpenStudy (anonymous):

i literally have no clue

OpenStudy (anonymous):

im sorry im stupid

OpenStudy (cwrw238):

x = square root of 9 = ?

OpenStudy (cwrw238):

what number multiplied by itself gives 9?

OpenStudy (anonymous):

obviously3

OpenStudy (anonymous):

how do i state the number and type of roots

OpenStudy (cwrw238):

right - then thats a root of the equation x^4 = 81 3^4 = 81 so x = 3 theres also another root x = -3 because (-3)^4 = 81

OpenStudy (cwrw238):

there are 2 roots (3,-3) which are real roots

OpenStudy (anonymous):

because with a square root a positive and negative of something by itself can equal the square root>?

OpenStudy (anonymous):

and theyre real because they arent imaginary, correct?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

times itself*

OpenStudy (anonymous):

ugh thank you

OpenStudy (cwrw238):

ok

OpenStudy (anonymous):

ill probably have more questions